OpenStudy (anonymous):

I need help solving this system with linear combination 2x-y=-6 5x+3y=-15

OpenStudy (anonymous):

Did you get the answer to the previous one?

OpenStudy (anonymous):

multiply first equation by 3

OpenStudy (anonymous):

yes @Aphrodite

OpenStudy (anonymous):

6x-3y=-18

OpenStudy (anonymous):

now add two equations together u will find that y cancels out

OpenStudy (anonymous):

@hahd can you please write out all the steps in one post? Its easier for me to understand that way

OpenStudy (anonymous):

multiply first equation by 3 6x-3y=-18 now add two equations together u will find that y cancels out 6x-3y=-6 + 5x+3y=-15 ------------- 11x+0=-21 x=-2